# Safe Interval Motion Planning for Quadrotors in Dynamic Environments

### v1

@@ IV-A1 Safe Intervals Generation for Edges
We represent the static environment with a 3D occupancy map and dynamic obstacles as bounded ellipsoids with trajectories parameterized by polynomials. For each edge in the graph, a cuboid covering it is generated and inflated by a margin equal to the Minkowski sum of the quadrotor and moving obstacles. Collision time stamps are determined by solving the points of intersection between the hyperplanes of the cube and moving obstacles’ trajectories. Finally Definition 2 is applied for calculating the safe intervals.
@@ IV-A2 Dynamic Connected Visibility Graph Construction
The algorithm for constructing the graph is detailed in Algorithm 1. Within a dynamic connected visibility graph, vertices are classified as either Guards or Connectors. Vertices are classified as Guard or Connectors depending on whether they can form edges with any other two vertices.
@@ V-B Spatial-Temporal Corridor Inflation
The spatial-temporal corridors generated around the initial paths only cover the edges of the geometric space, limiting the trajectory’s flexibility to further adjust around these paths. Hence, we need to inflate both the spatial and temporal corridors in 4-D space. Specifically, for each edge, we incrementally select two successive seed points along it, inflating the spatial corridors as cuboids in 3-D space while inflating the temporal corridors by checking potential collisions in these cuboids. Each spatial-temporal corridor is defined by a feasible temporal interval $(t_{l},t_{u})$ within which the cuboid is collision-free, and boundary points of the cuboid $\{b_{l},b_{u}\}$, where $b_{l}$ and $b_{u}$ are the lower bound and upper bound in each 3-D dimension, respectively. Fig. 4 provides a 1-D example of spatial-temporal corridor inflation.
@@ V-C Objectives Evaluation
We define the collision cost with dynamic obstacles as
@@
The spatial-temporal corridor cost is defined as the L1 norm of the distance between control points and corridors,
@@
We identify the cube corresponding to $\mathbf{Q}_{i}$ by locating the corridor whose temporal interval includes the knot of $\mathbf{Q}_{i}$.
Fig. 5 shows the 3-D simulation of back-end optimization with spatial-temporal corridors. The top-down view depicts a quadrotor’s trajectory from 0s to 9s. The initial B-spline trajectory is blue, corridors cubes are yellow, moving obstacles are in red while static obstacles are in black, each moving obstacle within the scope is attached with an id, and green cycles highlight the quadrotor’s position. The final trajectory with minimum control cost is represented in pink.
Figure 5: Simulation result (top-down view in 3-D environment) of the back-end optimization. Each cube is valid within its temporal intervals. The yellow cuboid demonstrates the spatial-temporal corridor.
@@ VI-B2 Trajectory Evaluations
To further validate the effectiveness of our framework, we evaluate our planner by comparing its performance with a baseline planner: TPRMO, which uses TPRM as the front-end method and optimizes it without the use of spatial-temporal corridors. The front-end and back-end planning framework is triggered when a collision is detected on an initial B-spline trajectory. We only evaluate trajectories that include valid front-end paths. For the baseline method, we include the Euclidean distances towards moving obstacles as cost functions and use the static obstacle avoidance strategy mentioned in [30]. We conducted 100 trials for each type of map. The results in Table I show that the success rate in one-time optimization (without re-optimizing upon failure), the average trajectory length, and duration are comparable between the two methods. However, the average control cost (integral of the squared jerk) is significantly reduced using our method. It demonstrates that our two-stage method, utilizing the spatial-temporal corridor, produces smoother trajectories. Re-parameterizing trajectory time within the corridor ensures both safety and efficiency.

### v2

@@ IV-A1 Safe Intervals Generation for Edges
We represent the static environment with a 3D occupancy map and dynamic obstacles as bounded ellipsoids with trajectories parameterized by polynomials. For each edge in the graph, a cuboid covering it is generated and inflated by a margin equal to the Minkowski sum of the quadrotor and moving obstacles. Collision time stamps are determined by solving the points of intersection between the hyperplanes of the cuboid and moving obstacles’ trajectories. Finally Definition 2 is applied for calculating the safe intervals.
@@ IV-A2 Dynamic Connected Visibility Graph Construction
The algorithm for constructing the graph is detailed in Algorithm 1. Within a dynamic connected visibility graph, vertices are classified as either Guards or Connectors. Vertices are classified as Connectors if they can connect with any other two Guards; otherwise, they are classified as Guards.
@@ V-B Spatial-Temporal Corridor Inflation
The front-end path is an initial spatial-temporal corridor, where edges define spatial corridors and time intervals set temporal corridors. However, the spatial corridors are confined to geometric edges, limiting trajectory flexibility. Hence, we need to inflate both the spatial and temporal corridors in 4-D space [25]. Specifically, for each edge, we incrementally select two successive seed points along it, inflating the spatial corridors as axis-aligned cuboids in 3-D space while inflating the temporal corridors by checking potential collisions in these cuboids. Each spatial-temporal corridor is defined by a feasible temporal interval $(t_{l},t_{u})$ within which the cuboid is collision-free, and boundary points of the cuboid $\{b_{l},b_{u}\}$, where $b_{l}$ and $b_{u}$ are the lower bound and upper bound in each 3-D dimension, respectively. Fig. 4 provides a 1-D example of spatial-temporal corridor inflation.
@@ V-C Objectives Evaluation
We also incorporate a collision cost for dynamic obstacles, as the generated trajectory may come too close to the boundaries of cuboids.
@@
The spatial-temporal corridor cost is defined as the L1 norm of the distance between control points and cuboids,
@@
We identify the cuboid corresponding to $\mathbf{Q}_{i}$ by locating the spatial corridor whose temporal interval includes the knot of $\mathbf{Q}_{i}$.
Fig. 5 visualizes back-end optimization with spatial-temporal corridors. The top-down view shows a quadrotor’s trajectory (0s–9s), the initial B-spline (blue), corridor cuboids (yellow), moving obstacles (red) with IDs, static obstacles (black), and the quadrotor’s position (green circles). The optimized trajectory with minimal control cost is in pink.
Figure 5: Simulation result (top-down view in 3-D environment) of the back-end optimization. Each cuboid is valid within its temporal intervals. The yellow cuboid demonstrates the spatial-temporal corridor.
@@ VI-B2 Trajectory Evaluations
To further validate the effectiveness of our framework, we evaluate our planner by comparing its performance with a baseline planner: TPRMO, which uses TPRM as the front-end method and optimizes it without the use of spatial-temporal corridors. The front-end and back-end planning framework is triggered when a collision is detected on an initial B-spline trajectory. Trajectories are generated and evaluated if the front-end planner can find a valid path. For the baseline method, we include the Euclidean distances towards moving obstacles as cost functions and use the static obstacle avoidance strategy mentioned in [30]. We conducted 100 trials for each type of map. The results in Table I show that the success rate in one-time optimization (without re-optimizing upon failure), the average trajectory length, and duration are comparable between the two methods. However, the average control cost (integral of the squared jerk) is significantly reduced using our method. It demonstrates that our two-stage method, utilizing the spatial-temporal corridor, produces smoother trajectories. Re-parameterizing trajectory time within the corridor ensures both safety and efficiency.